OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5

OpenStudy (solomonzelman):

Cuz it's -3x-9= -3x -3 -5 -3x+3= -3x -8 -9 ≠ -8 No Solution.

OpenStudy (solomonzelman):

-3x-9= -3x -3 -5 -9 ≠ -8

OpenStudy (anonymous):

thank you <3

OpenStudy (solomonzelman):

Anytime! ( I kept making mistakes in my work, and you still get the point...)

OpenStudy (anonymous):

solomon can you help me